OpenStudy (anonymous):

Simplify 6{[3(a-4)+17]-[3(2a-2)+5]}

OpenStudy (ybarrap):

\[36-18a\]

OpenStudy (anonymous):

This is what I got: 6{3(a-4) + 17] - [3(2a - 2) + 5]} = 6(3a - 12 + 17 - 6a - 6 + 5) = 6(3a + 5 - 6a - 1) = 6(-3a + 4) = -18a + 24

OpenStudy (anonymous):

6{[3(a-4)+17]-[3(2a-2)+5]} =6(3a-12+17-6a-6+5) =6(3a+5-6a-1) 6(8a-5a) =3a

OpenStudy (anonymous):

How did you get down to 3a? How did you remove the non-variables?